OpenStudy (anonymous):

20x^2 + 70x + 20 = 0

Parth (parthkohli):

Quadratic formula *whistles*

OpenStudy (anonymous):

Options are x = -7 +or- √33/4 x = 7 +or- √6/4 x = -70 +or-√33/4 or x = 70 +or-√6 /4

Parth (parthkohli):

ax^2 + bx + c...a is 20....b is 70...c is 20. Do you know the formula?

OpenStudy (anonymous):